# Retrograde predominance of small saturnian moons reiterates a recent retrograde collisional disruption

### v1

@@ 6.1 Simulated collision
Figure 7: An attempt to recreate the spread in orbit of the Mundilfari subgroup members (magenta) by simulating 2000 ejecta particles (green). The ejecta particles were created by adding $\Delta v$ = 200 m/s in a random direction to that of a chosen progenitor (black cross). The vertical bars through some the real moons indicate the range of orbital values that a moon oscillates over during a secular cycle, taken from Jacobson et al. (2022). The variations in $a$ are comparatively small and are thus not shown. The new moons do not have orbital elements ranges since they were announced after the study by Jacobson et al. (2022).
We generated 2000 ejecta particles by adding a fixed $\Delta{\bf v}$ in random directions to that of the progenitor velocity at a moment of collision, similar to what was done in Grav et al. (2003). To be able to realistically reproduce the Mundilfari subgroup, we have the criteria that at least one of the generated ejecta must overlap with the ranges of $a$, $e$, and $i$, found in Jacobson et al. (2022) of every moon in the Mundilfari subgroup. There are two caveats with this approach. Firstly, we did not look at the 3 dimensional $a-e-i$ space, just the $a-e$ and $a-i$ space. Secondly, when Jacobson et al. (2022) did their analysis, only one of the 64 new moons was announced, thus do not have computed orbital ranges. Thus, we cannot know whether the ejecta correctly recreate the full range of the now-larger Mundilfari subgroup members. We find that a minimum of 200 m/s is needed for $\Delta v$, occurring when the progenitor had $\Omega\simeq 0^{\circ}$, $\omega\simeq 270^{\circ}$, and $f\simeq 240^{\circ}$, to produce an ejecta field covering the ranges of almost all of the Mundilfari subgroup members (see Figure 7). It was much easier to fit the spread in $i$ compared to the spread in $e$ and $a$
6.2 Velocity Dispersion
@@
Turrini et al. (2008) used a velocity dispersion limit of 200 m/s to identify potential collisional families within the saturnian irregular moon population, where the velocity dispersion for a set of moons is obtained using Gauss equations. Although our minimum $\Delta v$ is the same as the velocity dispersion limit used by Turrini et al. (2008), the two velocities are not quite equivalent. Velocity dispersions are calculated using the mean orbital elements, whereas we fit the range of orbital elements. The Cluster C family defined in Table 3 of Turrini et al. (2008) contains 12 out 14 Mundilfari subgroup members known then, and additionally Griep and S/2006 S 1 (both of which have inclinations close to limits that define the Mundilfari subgroup). Due to the similarities in the members of the Mundilfari subgroup and the Cluster C family, the velocity dispersions of the two sets of moons should be very similar. Since Turrini et al. (2008) found that the Cluster C family has a velocity dispersion of 240 m/s, the velocity dispersion for the Mundilfari subgroup likely exceeds 200 m/s. If one accepts this criterion, the Mundilfari subgroup appears to be too dispersed to be a collisional family.
@@
Perhaps some members at the extrema $a$ and/or $e$ of the Mundilfari subgroup were not formed from the recent collision (thus causing the spread in $a$ and/or $e$ of the collisional family to appear larger than it actually is) or perhaps the orbital spread has been enlarged by subsequent collisions within the subgroup but not enough to grind down the steep size distribution. Alternately, it could be that there were two (or more) separate collisions that both occurred recently and which are slightly overlapping in orbital phase space. Li & Christou (2018) found that the gravitational interaction with Phoebe causes ejecta from the large moon to disperse over the age of the Solar System. However, the effect of Phoebe on the spread of the Mundilfari subgroup is likely minimal due to the potential recency of the subgroup and the lower rate of Phoebe interactions caused by large mutual inclinations.
@@
6.3 Colours
Numerous works have measured the colours of saturnian irregulars (Grav et al., 2003; Buratti et al., 2005; Grav & Bauer, 2007; Graykowski & Jewitt, 2018; Peña & Fuentes, 2022), although with the main focus of comparing the saturnian iregulars, along with irregular moons of other giant planets, to other populations of small bodies in the Solar System. There has been very little attempt to use colour information to identify potential collisional families. This is mainly due to the difficulty of getting high-quality colours of a large enough sample of saturnian irregulars, caused by the faintness of almost all of the moons coupled with their irregular shapes.
Interestingly, Grav & Bauer (2007) found that Mundilfari has a V-I colour of 0.52 $\pm$ 0.07, significantly bluer than almost any other saturnian irregular. The only saturnian irregular that has similar V-I is Phoebe with 0.64 $\pm$ 0.01. Thus far, V-I colours (or the equivalent g-i in the Sloan filter system) have not been measured for any other Mundilfari subgroup member. Aegir and Fornjot were included in the sample from Graykowski & Jewitt (2018), but they did not obtain observations in the I filter. Peña & Fuentes (2022) observed Skoll, Fornjot, and Loge but do not provide $g$-band magnitudes for those two moons. If the Mundilfari subgroup is indeed a collisional family, then we expect that other members will also exhibit blue V-I colours that Mundilfari is observed to have.
@@ 7 Conclusion
The Mundilfari subgroup, with differential size index $q$=6, has a significantly steeper size distribution slope compared to all other groups/subgroups. Due to the steep size distribution, we believe that the Mundilfari subgroup was created by the recent collision proposed by Ashton et al. (2021). A major concern, however, is the large orbital phase space that the Mundilfari subgroup occupies, which is larger than one might expect for the spread coming from the break-up speed of a collisional family. The recent collision hypothesis would be greatly strengthened if other members of the Mundilfari subgroup were found to have the very blue V-I colour of Mundilfari.
@@ 8 Acknowledgments
This work was supported by funding from the Natural Sciences and Engineering Research Council of Canada. Thanks to the CFHT Queue Observing team, especially Todd Burdullis and Daniel Devost, for helping us with the data acquisition process. We thank Stephen Gwyn for helpful discussions.

### v2

@@ 6.1 Gallic group
Figure 7: A thousand ejecta particles were simulated in attempt to recreate the spread in orbit of the Gallic group with S/2004 S 24 and without. The ejecta particles were created by adding $\Delta v$ = 80 m/s (red dots) or 330 m/s (blue dots) to that of the chosen progenitor (magenta cross). Time-averaged orbital elements from JPL of all of the moons in the subgroup are shown by black circles. The vertical bars through some the real moons indicate the range of orbital values that a moon oscillates over during a secular cycle, taken from Jacobson et al. (2022). The variations in $a$ are typically comparatively small and are thus not shown. The orbital elements ranges of the newest moons are not shown since they were announced after the Jacobson et al. (2022) study and are therefore not provided by JPL.
6.2 Mundilfari subgroup
As mentioned in section 5 there is evidence that the entire Mundilfari subgroup was created by a recent catastrophic collision. To test this, we used the approximate centre of the Mundilfari subgroup for $i$, $a$, and $e$ as the orbit of the progenitor, with values of 165, 0.126 au, and 0.28 respectively. We varied $\Omega$, $\omega$, and $f$ to find an ejecta field that best matched the orbits of the Mundilfari subgroup. Since this subgroup is more dispersed in orbital element space, we have increased the number of simulated ejecta to 2000.
Refer to captionRefer to caption
Figure 8: An attempt to recreate the spread in orbit of the Mundilfari subgroup members (black) by simulating 2000 ejecta particles (red). The ejecta particles were created by adding $\Delta v$ = 200 m/s in a random direction to that of a chosen progenitor (magenta cross). See Figure 7 caption for more details.
We find that a minimum of 200 m/s is needed for $\Delta v$ to produce an ejecta field that covers the ranges of almost all of the Mundilfari subgroup members (see Figure 8). This occurred when the progenitor had $\Omega\simeq 0^{\circ}$, $\omega\simeq 270^{\circ}$, and $f\simeq 240^{\circ}$. It was much easier to fit the spread in $i$ compared to the spread in $e$ and $a$.
6.2.1 Velocity Dispersion
@@
Based on hydrocode models, Nesvorný et al. (2003) set a velocity dispersion limit of 100 m/s to identify potential collisional families within the saturnian irregular moon population, where the velocity dispersion for a set of moons is obtained using Gauss equations. In Turrini et al. (2008), they decided to relax the limit to 200 m/s to account for the large oscillations that the orbits of irregular moons can have. Although our minimum $\Delta v$ is the same as the velocity dispersion limit used by Turrini et al. (2008), the two velocities are not quite equivalent. Velocity dispersions are calculated using the mean orbital elements, whereas we fit the range of orbital elements.
The Cluster C family defined in Table 3 of Turrini et al. (2008) contains 11 of 13 Mundilfari subgroup members known then, and additionally Griep, Farbauti, and S/2006 S 1 (all of which have inclinations close to the limits that define the Mundilfari subgroup, see section 5). Due to the similarities in the members of the Mundilfari subgroup and the Cluster C family, the velocity dispersions of the two sets of moons should be very similar. Since Turrini et al. (2008) found that the Cluster C family has a velocity dispersion of 240 m/s, the velocity dispersion for the Mundilfari subgroup likely exceeds 200 m/s. If one accepts this criterion, the Mundilfari subgroup appears to be too dispersed to be a collisional family.
@@
Perhaps some members at the extrema $a$ and/or $e$ of the Mundilfari subgroup were not formed from the recent collision (thus causing the spread in $a$ and/or $e$ of the collisional family to appear larger than it actually is) or perhaps the orbital spread has been enlarged by subsequent collisions within the subgroup but not enough to grind down the steep size distribution. Alternately, it could be that there were two (or more) separate collisions that both occurred recently and overlapped slightly in orbital phase space. Li & Christou (2018) found that the gravitational interaction with Phoebe causes ejecta from the large moon to disperse over the age of the Solar System. However, the effect of Phoebe on the spread of the Mundilfari subgroup is likely minimal due to the likely recency of the subgroup and, to a lesser extent, the lower rate of Phoebe interactions due to the larger mutual inclinations (than what Phoebe impact ejecta would have).
@@
6.2.2 Colours
Numerous works have measured the colours of saturnian irregulars (Grav et al., 2003; Buratti et al., 2005; Grav & Bauer, 2007; Graykowski & Jewitt, 2018; Peña & Fuentes, 2022), although with the main focus of comparing saturnian irregulars, along with irregular moons of other giant planets, to other populations of small bodies in the Solar System. There has been very little attempt to use colour information to identify potential collisional families. This is mainly due to the difficulty of getting high-quality colours of a large enough sample of saturnian irregulars, caused by the faintness of almost all of the moons coupled with their irregular shapes.
Interestingly, Grav & Bauer (2007) found that Mundilfari has a V-I colour of 0.52 $\pm$ 0.07, significantly bluer than almost any other saturnian irregular. The only saturnian irregular that has similar V-I is Phoebe with 0.64 $\pm$ 0.01. Thus far, V-I colours (or the equivalent g-i in the Sloan filter system) have not been measured for any other Mundilfari subgroup member. Aegir and Fornjot were included in the sample from Graykowski & Jewitt (2018), but they did not obtain observations in the I filter. Peña & Fuentes (2022) observed Skoll, Fornjot, and Loge but did not provide $g$-band magnitudes for these three moons. If the Mundilfari subgroup is indeed a collisional family, then we expect that other members will also exhibit blue V-I colours that Mundilfari is observed to have.
@@ 7 Conclusion
The Mundilfari subgroup, with differential size index qq=6, has a significantly steeper size distribution slope compared to all other groups/subgroups. Due to the steep size distribution, we believe that the Mundilfari subgroup was created by the recent collision proposed by Ashton et al. (2021). A major concern, however, is the large orbital phase space that the Mundilfari subgroup occupies, which is larger than one might expect for the spread coming from the break-up speed of a collisional family. The recent collision hypothesis would be greatly strengthened if other members of the Mundilfari subgroup were found to have the very blue V-I colour of Mundilfari.
@@ 8 Acknowledgments
This work was supported by funding from the Natural Sciences and Engineering Research Council of Canada. Thanks to the CFHT Queue Observing team, especially Todd Burdullis and Daniel Devost, for helping us with the data acquisition process. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency.